OpenStudy (anonymous):

factor the binomials completely. 4a^2 - 121b^2

OpenStudy (anonymous):

this is a difference of two squares which factors to the product of a sum and difference of two binomials; the formula is\[x^2-y^2=(x-y)(x+y)\]in your problem x=2a and y=11b\[4a^2-121b^2\]\[=(2a)^2-(11b)^2\]\[=(2a-11b)(2a+11b)\]

OpenStudy (anonymous):

perfect, ty again

OpenStudy (anonymous):

ur welcome :})